OpenStudy (anonymous):

how can i change 7256 to a fraction?

OpenStudy (unklerhaukus):

\[7256/1\]\[=3628/2\]\[=907/8 \]

OpenStudy (anonymous):

how did u do that?

OpenStudy (unklerhaukus):

well any number can be converted to a fraction by dividing by one, looking at the numerator and see it that it is even means it can be divide by two, do this a few times

OpenStudy (unklerhaukus):

907 is a prime number and cannot be broken down any further

OpenStudy (anonymous):

oh okay i see! :) thank you!!!

OpenStudy (unklerhaukus):

(by the way i had to look up a table of prime numbers to know that 907 is prime)

OpenStudy (anonymous):

okay yeah i did not know that either ha